# An exploration of low crossing and chiral cosmetic bands with grid diagrams

### v2

@@ Abstract.
[…]
@@ 2. Grid diagrams and band attachments
A grid diagram of size $n$ is a $n\times n$ square grid, divided into $n^{2}$ small squares. Two sets of $n$ markings, usually denoted by $X$ and $O$, are placed within the grid, so that no two markings of the same kind are in the same row or column, and each small square contains at most one marking (see Figure 1).
Figure 1. A grid diagram of size $9$, representing the knot $7_{3}$.
@@
Figure 2. (a) The top row shows that the link diagram in the top left changes by a planar isotopy if both pairs of markings are moved. The mid row shows that swapping markings of the same type results in a coherent band attachment. Note that the two possibilities produce equivalent diagrams. The bottom row shows that swapping two markings of different types results in a non-coherent band attachment. Note that in this case the two possibilities produce non-equivalent links. (b) The two moves on the top diagram correspond to non-coherent band attachments yielding equivalent links.
@@
A band attachment is called coherent if the two links it relates have a different number of components, and non-coherent otherwise.
@@
We say that two consecutive rows are interleaved if the arcs spanned by the markings partially overlap (see the top diagram in Figure 2(b)), and non-interleaved otherwise.
We can non-trivially change the markings in these two rows in five possible ways, while still producing a valid grid diagram. The first possibility is to move one pair of markings upwards and the other pair downards. The effect on the link is either a crossing change or equivalent to a planar isotopy on the underlying knot diagram. These two alternatives depend on whether the two markings were interleaved or not, see the top row of Figure 2(a) for the non-interleaving case. The second possibility is to swap two markings among the two rows and leave the other two in the same position. In this case the effect of the move depends on whether we move two markings of the same kind (e.g. two $O$s), or not. In the first case, the move corresponds to a coherent band attachment, and in the second case to a non-coherent one, see the mid and bottom row in Figure 2(a). Depending on whether the rows/columns are interleaved or not, and whether the corresponding arcs are oriented in the same direction or not, exchanging one pair $X\leftrightarrow O$ among the two rows can yield non-equivalent band attachments, see Figure 2(a) mid and bottom row, and Figure 2(b) for an example. Note that this can occur in the coherent and non-coherent case alike.
@@ 3. Methods and results
Our computational exploration was achieved by using software in [15], based on GridPyM [1], see Software and data availability.
@@
For every chiral representative (up to orientation) of the knot types considered, we select grid diagrams minimising the grid number. We then randomise each of these seed configurations by applying a customisable amount of random grid moves. This is achieved by using the scramble_grid function on GridPyM. In a standard run, we apply $1000$ different moves a total of $400$ times for each seed configuration, yielding a collection of complex and high-grid number diagrams, see Figure 3. On average, this results in performing over $32,000$ band attachments per seed configuration, per run. The number of runs we performed varied depending on the knot type.
Once this collection is created, we perform all possible non-coherent band attachments in each generated diagram, using the built-in methodology explained in Section 2. Sage’s get_knotinfo() functionality [27] is then used to determine the resulting knot types (following KnotInfo’s conventions [19]). More details are available in the code repository [15], see also Software and data availability.
@@
Figure 4. Top: A band attachment from $7_{4}$ to $3_{1}m\#3_{1}m$. Bottom: an explicit example of the band attachment realised in the grids model found by our computations. The band (in red) is attached on row 28.
@@ 3.1. The $\mathrm{H}(2)$-adjacency tables
We collect a summary of our results in three tables: 2, 3, 4, containing prime knots with $\leq 7$ crossings, prime knots with $8$ crossings and composite knots with $\leq 8$ crossings respectively. Overall, we find a total of 33 previously unknown distance one pairs of knots with fewer than $9$ crossings. Two of these pairs connect knots with at most $7$ crossings: $7_{4}$ with $3_{1}m\#3_{1}m$, and $7_{3}$ with $7_{3}m$. Note that we are disregarding orientation in our computations.
@@
Examples of specific band attachments relating each distance one pair (excluding cosmetic ones) are available online [15], see Software and data availability.
@@ 4. Chirally cosmetic band attachments
We conclude with a table of the relative occurrences of chiral cosmetic band attachments in the grid model. It is noteworthy to point out that the frequency of occurences is between $1$ (for $8_{20}$) and $3$ (for $8_{8}$) orders of magnitude higher than in the analogous analysis by Moore and Vazquez [21]. Note that this phenomon is coherent with the fact that the grid diagram model is overly knotted (i.e. knots with high crossing number appear at low grid size) compared to other models.
@@
Table 1. Probability of occurences for chirally cosmetic band attachments in our data set cf[21, Table 1]
Software and data availability
The code for our computational exploration can be accessed in CK’s GitHub repository [15]. The software is based based on GridPyM [1]. The repository also contains one example for each band attachment found.
@@
We consider a knot diagram as shown in Figure 7. The rectangular box $R$ represents a 2-string tangle inserted there. We assume that the tangle $R$ admits a reflection symmetry along the vertical axis; that is, $R\cong\reflectbox{$R$}$.

### v3

@@ Abstract.
We computationally explore non-coherent band attachments between low crossing number knots, using grid diagrams. We significantly improve the current $\mathrm{H}(2)$-distance table. In particular, we find two new distance one pairs with at most seven crossings: one between $3_{1}\#3_{1}$ and $7_{4}m$, and a chirally cosmetic one for $7_{3}$. We further determine a total of 33 previously unknown $\mathrm{H}(2)$-distance one pairs for knots with up to $8$ crossings. The appendix by Kazuhiro Ichihara, In Dae Jong and Masakazu Teragaito contains a construction explaining the existence of chirally cosmetic bands for an infinite family of knots, including $5_{1},\,7_{3}$ and $8_{8}$.
@@ 2. Grid diagrams and band attachments
A grid diagram of size $n$ is an $n\times n$ square grid, divided into $n^{2}$ small squares. Two sets of $n$ markings, usually denoted by $X$ and $O$, are placed within the grid, so that no two markings of the same kind are in the same row or column, and each small square contains at most one marking (see Figure 1).
Figure 1. A grid diagram of size 99, representing the knot $7_{3}$.
@@
Figure 2. (a) Depicts a single portion of a grid diagram (on the left) being changed in each of the five possible ways (on the right). The top row shows that the link diagram on the left changes by a planar isotopy if both pairs of markings are moved. The mid row shows that swapping markings of the same type in the left diagram results in a coherent band attachment. Note that the two possibilities produce equivalent diagrams. The bottom row shows that swapping two markings of different types in the left diagram results in a non-coherent band attachment. Note that in this case the two possibilities might produce non-equivalent links. (b) The two moves on the left diagram correspond to non-coherent band attachments yielding equivalent links.
@@
A band attachment is called coherent if the two links it relates have different numbers of components, and non-coherent otherwise.
@@
We say that two consecutive rows are interleaved if the arcs spanned by the markings partially overlap (see the left diagram in Figure 2(b)), and non-interleaved otherwise.
We can non-trivially move pairs of markings vertically within the two rows in five possible ways, while still producing a valid grid diagram. The first possibility is to move one connected XO pair of markings upwards and the other pair downwards. The effect on the link is either a crossing change or equivalent to a planar isotopy on the underlying knot diagram. These two alternatives depend on whether the two markings were interleaved or not, see the top row of Figure 2(a) for the non-interleaving case. The second possibility is to move a single marking upward and another non-connected single marking downward and leave the other two in the same position. In this case the effect of the move depends on whether we move two markings of the same kind (e.g. two $O$s), or not. In the first case, the move corresponds to a coherent band attachment, and in the second case to a non-coherent one, see the mid and bottom row in Figure 2(a). Note that in the latter case the result is an invalid grid $G$ as the link orientation is not preserved by the move. However, it is possible to exhibit a valid grid $G^{\prime}$ whose unlabelled markings are the same as $G$. In our computations, non-coherent band attachments are achieved using built-in functions in GridPym, see [1] and Section 5 for further details and documentation. Depending on whether the rows/columns are interleaved or not, and whether the corresponding arcs are oriented in the same direction or not, exchanging one pair $X\leftrightarrow O$ among the two rows can yield non-equivalent band attachments, see Figure 2(a) mid and bottom row, and Figure 2(b) for an example. Note that this can occur in the coherent and non-coherent case alike.
@@ 3. Methods and results
Our computational exploration was achieved by using software in [15], based on GridPyM [1], see Software and data availability (Section 5).
@@
For every chiral representative (up to orientation) of the knot types considered, we select grid diagrams minimising the grid number. We then randomise each of these seed configurations by applying a customisable amount of random grid moves. This is achieved by using the scramble_grid function on GridPyM. This function takes as input a grid diagram $G$ and a positive integer $m$. The output is a new grid diagram, representing the same knot type as $G$, obtained by applying a sequence of $m$ random Cromwell moves. For a detailed description of the function and for GridPym’s documentation see [1].
In a standard run, for each seed configuration $G$, we compute $400$ different randomisations by applying the scramble_grid function with $m=1000$. This yields a collection of $400$ complex and high-grid number diagrams, see Figure 3. Once this collection is created, we search for all the possible non-coherent band attachments in each of the $400$ generated diagrams. We do this by following the built-in methodology explained in Section 2, using GridPym’s functions, see Section 5, and [15, 1] for a complete description and further documentation. On average, this results in performing over $32,000$ band attachments per seed configuration, per run. The number of runs we performed varied depending on the knot type.
Sage’s get_knotinfo() functionality [27] is then used to determine the resulting knot types (following KnotInfo’s conventions [19]). More details are available in the code repository [15], see also Section 5.
@@
Figure 4. Top: A non-coherent band attachment from $7_{4}$ to $3_{1}m\#3_{1}m$. Bottom: an explicit example of the band attachment realised in the grids model found by our computations. The band (in red) is attached on row 28.
@@ 3.1. The $\mathrm{H}(2)$-adjacency tables
We collect a summary of our results in three tables: Tables 2, 3, 4, containing prime knots with $\leq 7$ crossings, prime knots with $8$ crossings and composite knots with $\leq 8$ crossings respectively. Overall, we find a total of 33 previously unknown distance one pairs of knots with fewer than $9$ crossings. Two of these pairs connect knots with at most $7$ crossings: $7_{4}$ with $3_{1}m\#3_{1}m$ (see Figure 4), and $7_{3}$ with $7_{3}m$. Note that we are disregarding orientation in our computations.
@@
Examples of specific band attachments relating each distance one pair (excluding cosmetic ones) are available online [15], see the Software and data availability, Section 5.
@@ 4. Chirally cosmetic band attachments
We conclude with a table (Table 1) of the relative occurrences of chiral cosmetic band attachments in the grid model. It is noteworthy to point out that the frequency of occurences is between $1$ (for $8_{20}$) and $3$ (for $8_{8}$) orders of magnitude higher than in the analogous analysis by Moore and Vazquez [21]. Note that this phenomenon is coherent with the fact that the grid diagram model is overly knotted (i.e. knots with high crossing number appear at low grid size) compared to other models.
@@
Table 1. Probability of occurences for chirally cosmetic band attachments in our data set cf. [21, Table 1]
5. Software and data availability
The code for our computational exploration can be accessed in CK’s GitHub repository [15]. The software is based on GridPyM [1]. The repository also contains one example for each band attachment found.
@@ 5. Software and data availability
We consider a knot diagram as shown in Figure 7. The rectangular box $R$ represents a 2-string tangle inserted there, and $R^{*}$ denotes the mirror image of the tangle $R$. We assume that the tangle $R$ admits a reflection symmetry along the vertical axis; that is, $R\cong\reflectbox{$R$}$.